# ReNF: Rethinking the Design Space of Neural Long-Term Time Series Forecasters

### v3

@@ Abstract
Neural Forecasters (NFs) are a cornerstone of Long-term Time Series Forecasting (LTSF). However, progress has been hampered by an overemphasis on architectural complexity at the expense of fundamental forecasting principles. In this work, we return to first principles to redesign the LTSF paradigm. We begin by introducing a Multiple Neural Forecasting Principle that provides a theoretical motivation for our approach. We propose Boosted Direct Output (BDO), a novel forecasting paradigm that synergistically hybridizes the causal nature of Auto-Regressive (AR) models with the stability of Direct Output (DO). In addition, we stabilize the learning process by smoothly tracking the model’s parameters. Extensive experiments show that these principled improvements enable a simple MLP to achieve state-of-the-art performance, outperforming recent, complex models in nearly all cases, without any specific considerations in the area. Finally, we empirically verify our principle, establishing a dynamic performance bound and identifying promising directions for future research. The code for review is available at: .
@@ 1 Introduction
The progression of any real-world event is a unique, non-repeatable process, often governed by chaotic dynamics we cannot perfectly measure (Robertson, 1929) or describe. This implies that any observed time series is just one stochastic sample from a complex underlying system. Consequently, a fundamental open question persists: how can we best estimate the long-term future states depending only on a single, observed historical sequence?
Deep Neural Networks (DNNs) have recently received increasing attention in Long-Term Time Series Forecasting (LTSF) (Kim et al., 2025). Their ability to model high-dimensional, non-linear dependencies makes the Neural Forecaster (NF) a promising tool for capturing complex temporal dynamics and implicit dependencies (Laiz et al., 2024). However, the literature reveals several critical problems that should be addressed to unlock the full potential of DNNs in this domain.
One confusing puzzle is that both advanced architectures, like transformer-based (Nie et al., 2023) and simple linear-based models (Zeng et al., 2023), are reported to achieve state-of-the-art performance interchangeably while their complexity varies significantly. This is partly because some datasets prefer parsimonious (Deng et al., 2024) while the others have sufficient volume for complex models to fit. But perhaps a deeper issue is the insufficient exploration of the networks’ intrinsic capabilities. As indicated by (Lu et al., 2025), many NFs contain redundant components and are used selectively depending on the property of the data, resulting in their full potential remaining untapped without extensive tuning or regularizations.
Furthermore, the field’s focus has shifted towards designing specialized modules for properties that are believed to be beneficial for LTSF, such as multi-scale (Wang et al., 2024b) and non-stationary (Liu et al., 2022). However, the progress has become erratic because these architectural additions often yield subtle gains while overlooking fundamental principles. We re-emphasize that progress in this field becomes circuitous when we introduce advanced models merely according to the empirical conclusions from other realms, but overlook the detailed instructions for their full utilization. A primary and more direct path to advancement may lie in fundamentally improving the training stability and generalization capabilities of NFs themselves.
@@
  • We propose the Multiple Neural Forecasting Principle (MNFP) with empirical evaluations, which provides a formal motivation for employing multiple Neural Forecasters (NFs) in Long-Term Time Series Forecasting (LTSF).
@@ 2.2 Theoretical Motivation for Multiple Forecasts
  • (A3) Each element in an output series is generated independently.
@@
We emphasize that condition (A4) distinguishes between the observed real series and the expected true series. This frames the observed data not as the absolute ground truth, but as a single, near-optimal sample from the underlying stochastic process. While it is feasible to theoretically approach the expected forecast, the empirical gap between any observation $x_{t}$ and the true expectation $\mu_{t}$ is practically irreducible.
The proof can be found in Appendix C. To interpret, the shown bound value provides us with some intuitions: for instance, the forecasting error of the NFM naturally grows with the forecast horizon $T$ and the data range $\lambda$, because of the expanded solution space.
The most critical insight of the proposition is that an accurate forecast is theoretically achievable even with a rather weak generator, provided a sufficiently large number of candidate outputs is produced. This principle is rooted in the well-established theory of ensemble methods (Bates & Granger, 1969), which demonstrates that combining multiple diverse estimators can lead to a more accurate and robust final estimate. Furthermore, higher-quality historical data and a more powerful NFM (which reduces the estimation error $b$) both serve to constrain the solution space, leading to better predictions.
@@
In our experiments, we will empirically show the potential of this combination on real-world datasets. In practice, however, directly creating an optimal combination function without access to future information is a significant challenge. Therefore, designing a framework that can implicitly and effectively leverage these multiple forecasts is one of the key principles guiding our subsequent work.
@@ 2.3 Novel Forecasting Paradigm
The current frameworks in LTSF involve two main strategies: 1-step Auto Regressive (AR(1)) and Direct Output (DO). While AR models perform consistently with recurrent state space models like RNNs (Siami-Namini et al., 2019), they are known to suffer from significant error accumulation in long-horizon forecasting and are empirically evidenced to largely fall behind DO in LTSF (Zhou et al., 2021), making a strong reason for the dominant position of DO in LTSF. However, the shortcomings of the DO approach itself have been largely overlooked. We argue that despite its simplicity, the most significant weakness of DO, particularly when compared to AR, is its lack of inherent causality.
@@ 2.4 Model Architecture
We construct our NFs using only MLP and linear layers for three primary reasons: 1) As foundational deep learning modules, improvements demonstrated on them are broadly applicable and convincing. 2) Their computational efficiency (low FLOPs) facilitates rapid experimentation and verification. 3) Recent work has shown that simple MLP-based architectures are often sufficient for a wide range of forecasting tasks (Ekambaram et al., 2023).
@@
We employ RevIN (Kim et al., 2021) as the pre-normalization for the initial input data to reduce the distribution discrepancy between the training and evaluation phases. For consistency, we also apply pre-normalizations to the input of each sub-forecaster. Additionally, we apply dropout before the initial linear projection to prevent the model from being overly dependent on the history observation and the concatenated information in the data space.
@@ 2.5 Learning Objective
To train a deterministic NF, we adopt the hybrid loss function as used in (Liu et al., 2025a), which is a convex combination of the Mean Absolute Error (MAE) in both the time and frequency domains. The frequency-domain component has been shown to be effective at reducing spurious autocorrelations in the labels (Wang et al., 2024a).
@@ 2.6 Smoothing the learning for time series
Notably, training on standard benchmark datasets is often unstable, partly due to sparse or redundant information. The non-stationary nature of real-world time series, combined with chronological data splitting, leads to both internal (batch-to-batch) and external (train/validation/test) distribution mismatches. This causes the optimization path to be heterogeneous, rendering the learning process for NFs ineffective. Specifically, internal data redundancy can cause the model to overfit repeated or spurious patterns (Liu et al., 2025a), harming generalization, while external distribution shift creates inconsistencies across the different phases of training. While normalization techniques like RevIN address the train-test shift, a critical problem remains between validation and testing.
@@
To mitigate these unexpected effects, we propose smoothing the training trajectory by employing a ”shadow” model for evaluation. We achieve this efficiently using an Exponential Moving Average (EMA) to track the parameters of the online model, a technique proven effective in self-supervised learning in self-supervised learning (LeCun, 2022) and generative models (Song et al., 2020). Specifically, let $\theta$ be the parameters of the online model being trained, and $\theta^{\prime}$ be the parameters of the shadow model. Then the shadow model’s parameters are upgraded as follows at each iteration.
@@ 3 Experiment
Baselines and Datasets. We compare our model with a suite of recent state-of-the-art methods, including TimeBridge (Liu et al., 2025a), DUET (Qiu et al., 2025), TimeDistill (Ni et al., 2025), Timer-XL (Liu et al., 2025b), iTransformer (Liu et al., 2024), TimeMixer (Wang et al., 2024b), PatchTST (Nie et al., 2023), Crossformer (Zhang & Yan, 2023), and Dlinear (Zeng et al., 2023). We use widely used benchmark datasets in this area, including electricity (ETTh1, ETTh2, ETTm1, ETTm2, Electricity), environment (Weather), energy (Solar-Energy), and transportation (Traffic). Detailed descriptions of these datasets are provided in Appendix B.
Setups. All experiments were conducted on a single NVIDIA 4090 GPU with 24GB of memory, using the Adam optimizer (Kingma & Ba, 2014) and a fixed random seed of 2021 for reproducibility. Results for all baseline models were reproduced using their official source code and optimal configurations. For our model, ReNF, we searched for the optimal learning rate in the range from 0.0001 to 0.005, the EMA decay rates in the range from 0.99 to 0.999, and the number of layers from 2 to 8. We apply ReNF-$\alpha$ to ETTh1 and ETTh2 datasets, and ReNF-$\beta$ to others. The look-back window of ReNF is searched over $\{336,512,720\}$ for the best performance. The drop_last bug is corrected following the TFB benchmark (Qiu et al., 2025).
@@ Appendix A Related Work
We note that ReNF is not the first work to address LTSF using multiple short-term forecasts. A pioneering and highly relevant study by (Nguyen & Chan, 2004) proposed a machine called MNN, which generates long-term forecasts using multiple neural networks, each responsible for a specific interval of the output. The essential distinction between MNN and our approach is that we generate sub-forecasts recursively, with the explicit goal of injecting causality into the DO strategy. In contrast, each sub-network in MNN functions as a standalone AR(1) model. This difference stems from our distinct motivation: whereas MNN was designed primarily to mitigate the error accumulation of one-step-ahead NN models, our work begins with a foundational forecasting proposition. Based on this proposition, we leverage modern deep learning techniques to effectively utilize multiple sub-forecasts, achieving strong empirical results on a wide range of real-world datasets, which constitutes one of our main contributions.
A recently proposed work using the multiple-choice learning paradigm (Cortés et al., 2025) is also relevant, as it demonstrates an ability to handle the diverse and multi-modal nature of the future from a probabilistic perspective. Furthermore, the characterization of TimeMCL as a conditional stationary quantizer for time series may offer additional theoretical support and interpretations for our framework.
Our work also connects to classic ensemble methods in machine learning (Mohri et al., 2018). For instance, the recursive workflow of ReNF is analogous to gradient-boosted regression models (Chen & Guestrin, 2016), which construct a strong predictor from multiple weak ones. Additionally, the collaborative training can be viewed as a form of bagging (Breiman, 1996), which effectively resamples the data to train diverse forecasters. In this context, an even more recent study (Modi & Pan, 2025) re-certified the benefits of such ensemble methods for enhancing Transformer-based NFs, providing further empirical support for the direction of our research.
Forecast Combinations. Forecast combination is a classic technique for improving forecast accuracy and robustness by leveraging the diverse strengths of multiple models (Clemen, 1989). In this work, we extend this concept (Bates & Granger, 1969) to the domain of deep learning for long-term time series forecasting. Rather than combining distinct, parallel forecasters into a hybrid model (Zhang, 2003), our framework achieves this goal efficiently within a single, structured approach for generating and implicitly combining forecasts within a single neural network. By recursively stacking sub-forecasts, similar in spirit to N-BEATS (Oreshkin et al., 2019), we ensure that the representation learning is both deep and diverse, allowing the LSTF task to benefit from the full power of modern machine learning.
@@ Appendix C Proof of the approximate proposition
  • (A3) Each element in an output series is generated independently. (This can be satisfied in the DO framework as discussed in Sec.2.3.)
@@
$|\mu_{t}-\frac{1}{c}\sum_{i=1}^{c}{\tilde{y}_{t}^{(i)}}|\leq|\hat{\mu}_{t}-\frac{1}{c}\sum_{i=1}^{c}{\tilde{y}_{t}^{(i)}}|+|\hat{\mu}_{t}-\mu_{t}|\leq|\hat{\mu}_{t}-\frac{1}{c}\sum_{i=1}^{c}{\tilde{y}_{t}^{(i)}}|+b.~~(\text{A4})$ (15)
@@
Note that the above proof is based on a univariate time series; however, it is easy to extend the result to the multivariate case using a similar process in a certain normed vector space.

### v4

@@ Abstract
Neural Forecasters (NFs) are a cornerstone of Long-term Time Series Forecasting (LTSF). However, progress has been hampered by an overemphasis on architectural complexity at the expense of fundamental forecasting principles. In this work, we return to first principles to redesign the LTSF paradigm. We begin by introducing a Multiple Neural Forecasting Proposition that provides a theoretical motivation for our approach. We propose Boosted Direct Output (BDO), a novel forecasting paradigm that synergistically hybridizes the causal nature of Auto-Regressive (AR) models with the stability of Direct Output (DO). In addition, we stabilize the learning process by smoothly tracking the model’s parameters. Extensive experiments show that these principled improvements enable a simple MLP to achieve state-of-the-art performance, outperforming recent, complex models in nearly all cases, without any specific considerations in the area. Finally, we empirically verify our proposition, establishing a dynamic performance bound and identifying promising directions for future research. The code for review is available at: .
@@ 1 Introduction
The progression of any real-world event is a unique, non-repeatable process, often governed by chaotic dynamics we cannot perfectly measure (robertson1929uncertainty) or describe. This implies that any observed time series is just one stochastic sample from a complex underlying system. Consequently, a fundamental open question persists: how can we best estimate the long-term future states depending only on a single, observed historical sequence?
Deep Neural Networks (DNNs) have recently received increasing attention in Long-Term Time Series Forecasting (LTSF) (survey2025). Their ability to model high-dimensional, non-linear dependencies makes the Neural Forecaster (NF) a promising tool for capturing complex temporal dynamics and implicit dependencies (selfcontrastivenonlinear). However, the literature reveals several critical problems that should be addressed to unlock the full potential of DNNs in this domain.
One confusing puzzle is that both advanced architectures, like transformer-based (patchtst) and simple linear-based models (dlinear), are reported to achieve state-of-the-art performance interchangeably while their complexity varies significantly. This is partly because some datasets prefer parsimonious (deng2024parsimony) while the others have sufficient volume for complex models to fit. But perhaps a deeper issue is the insufficient exploration of the networks’ intrinsic capabilities. As indicated by (lu2025timecapsule), many NFs contain redundant components and are used selectively depending on the property of the data, resulting in their full potential remaining untapped without extensive tuning or regularizations.
Furthermore, the field’s focus has shifted towards designing specialized modules for properties that are believed to be beneficial for LTSF, such as multi-scale (wang2024timemixer) and non-stationary (nonstarionarytransformer). However, the progress has become erratic because these architectural additions often yield subtle gains while overlooking fundamental principles. We re-emphasize that progress in this field becomes circuitous when we introduce advanced models merely according to the empirical conclusions from other realms, but overlook the detailed instructions for their full utilization. A primary and more direct path to advancement may lie in fundamentally improving the training stability and generalization capabilities of NFs themselves.
@@
  • We propose the Multiple Neural Forecasting Proposition (MNFP) with empirical evaluations, which provides a formal motivation for employing multiple Neural Forecasters (NFs) in Long-Term Time Series Forecasting (LTSF).
@@ 2.2 Theoretical Motivation for Multiple Forecasts
The proof and further analysis can be found in Appendix C. To interpret, the shown bound value provides us with some intuitions: for instance, the forecasting error of the NFM naturally grows with the forecast horizon $T$ and the data range $\lambda$, because of the expanded solution space. We emphasize that condition A3 distinguishes between the observed real series and the expected true series. This frames the observed data not as the absolute ground truth, but as a single, near-optimal sample from the underlying stochastic process. While it is feasible to theoretically approach the expected forecast, the empirical gap between any observation $x_{t}$ and the true expectation $\mu_{t}$ is practically irreducible.
The most critical insight of the proposition is that an accurate forecast is theoretically achievable even with a rather weak generator, provided a sufficiently large number of candidate outputs is produced. This proposition is rooted in the well-established theory of ensemble methods (bates1969combination), which demonstrates that combining multiple diverse estimators can lead to a more accurate and robust final estimate. Furthermore, higher-quality historical data and a more powerful NFM (which reduces the estimation error $b$) both serve to constrain the solution space, leading to better predictions. Inspired by this, we can propose the following conception,
@@
In our experiments, we will empirically show the potential of this combination on real-world datasets. In practice, however, directly creating an optimal combination function without access to future information is a significant challenge (combiningreview). Therefore, designing a framework that can implicitly and effectively leverage these multiple forecasts is one of the key principles guiding our subsequent work.
@@ 2.3 Novel Forecasting Paradigm
The current frameworks in LTSF involve two main strategies: 1-step Auto Regressive (AR(1)) and Direct Output (DO). While AR models perform consistently with recurrent state space models like RNNs (lstm4ltsf), they are known to suffer from significant error accumulation in long-horizon forecasting and are empirically evidenced to largely fall behind DO in LTSF (zhou2021informer), making a strong reason for the dominant position of DO in LTSF. However, the shortcomings of the DO approach itself have been largely overlooked. We argue that despite its simplicity, the most significant weakness of DO, particularly when compared to AR, is its lack of inherent causality.
@@ 2.4 Model Architecture
We construct our NFs using only MLP and linear layers for three primary reasons: 1) As foundational deep learning modules, improvements demonstrated on them are broadly applicable and convincing. 2) Their computational efficiency (low FLOPs) facilitates rapid experimentation and verification. 3) Recent work has shown that simple MLP-based architectures are often sufficient for a wide range of forecasting tasks (ekambaram2023tsmixer).
@@
We employ RevIN (kim2021reversible) as the pre-normalization for the initial input data to reduce the distribution discrepancy between the training and evaluation phases. For consistency, we also apply pre-normalizations to the input of each sub-forecaster. Additionally, we apply dropout before the initial linear projection to prevent the model from being overly dependent on the history observation and the concatenated information in the data space.
@@ 2.5 Learning Objective
To train a deterministic NF, we adopt the hybrid loss function as used in (liu2024timebridge), which is a convex combination of the Mean Absolute Error (MAE) in both the time and frequency domains. The frequency-domain component has been shown to be effective at reducing spurious autocorrelations in the labels (wang2024fredf).
@@ 2.6 Smoothing the learning for time series
Notably, training on standard benchmark datasets is often unstable, partly due to sparse or redundant information. The non-stationary nature of real-world time series, combined with chronological data splitting, leads to both internal (batch-to-batch) and external (train/validation/test) distribution mismatches. This causes the optimization path to be heterogeneous, rendering the learning process for NFs ineffective. Specifically, internal data redundancy can cause the model to overfit repeated or spurious patterns (liu2024timebridge), harming generalization, while external distribution shift creates inconsistencies across the different phases of training. While normalization techniques like RevIN address the train-test shift, a critical problem remains between validation and testing.
@@
To mitigate these unexpected effects, we propose smoothing the training trajectory by employing a ”shadow” model for evaluation. We achieve this efficiently using an Exponential Moving Average (EMA) to track the parameters of the online model, a technique proven effective in self-supervised learning in self-supervised learning (jepa) and generative models (song2020scoreSDE). Specifically, let $\theta$ be the parameters of the online model being trained, and $\theta^{\prime}$ be the parameters of the shadow model. Then the shadow model’s parameters are upgraded as follows at each iteration.
@@ 3 Experiment
Baselines and Datasets. We compare our model with a suite of recent state-of-the-art methods, including TimeBridge (liu2024timebridge), DUET (qiu2025duet), TimeDistill (ni2025timedistill), Timer-XL (timerxl), iTransformer (liu2023itransformer), TimeMixer (wang2024timemixer), PatchTST (patchtst), Crossformer (zhang2023crossformer), and Dlinear (dlinear). We use widely used benchmark datasets in this area, including electricity (ETTh1, ETTh2, ETTm1, ETTm2, Electricity), environment (Weather), energy (Solar-Energy), and transportation (Traffic). Detailed descriptions of these datasets are provided in Appendix B.
Setups. All experiments were conducted on a single NVIDIA 4090 GPU with 24GB of memory, using the Adam optimizer (kingma2014adam) and a fixed random seed of 2021 for reproducibility. Results for all baseline models were reproduced using their official source code and optimal configurations. For our model, ReNF, we searched for the optimal learning rate in the range from 0.0001 to 0.005, the EMA decay rates in the range from 0.99 to 0.999, and the number of layers from 2 to 8. We apply ReNF-$\alpha$ to ETTh1 and ETTh2 datasets, and ReNF-$\beta$ to others. The look-back window of ReNF is searched over $\{336,512,720\}$ for the best performance. The drop_last bug is corrected following the TFB benchmark (qiu2025duet).
@@ Appendix A Related Work
We note that ReNF is not the first work to address LTSF using multiple short-term forecasts. A pioneering and highly relevant study by (nguyen2004multiple) proposed a machine called MNN, which generates long-term forecasts using multiple neural networks, each responsible for a specific interval of the output. The essential distinction between MNN and our approach is that we generate sub-forecasts recursively, with the explicit goal of injecting causality into the DO strategy. In contrast, each sub-network in MNN functions as a standalone AR(1) model. This difference stems from our distinct motivation: whereas MNN was designed primarily to mitigate the error accumulation of one-step-ahead NN models, our work begins with a foundational forecasting proposition. Based on this proposition, we leverage modern deep learning techniques to effectively utilize multiple sub-forecasts, achieving strong empirical results on a wide range of real-world datasets, which constitutes one of our main contributions.
A recently proposed work using the multiple-choice learning paradigm (timemcl) is also relevant, as it demonstrates an ability to handle the diverse and multi-modal nature of the future from a probabilistic perspective. Furthermore, the characterization of TimeMCL as a conditional stationary quantizer for time series may offer additional theoretical support and interpretations for our framework.
Our work also connects to classic ensemble methods in machine learning (mohri2018foundationsofml). For instance, the recursive workflow of ReNF is analogous to gradient-boosted regression models (chen2016xgboost), which construct a strong predictor from multiple weak ones. Additionally, the collaborative training can be viewed as a form of bagging (breiman1996bagging), which effectively resamples the data to train diverse forecasters. In this context, an even more recent study (emsenble4trans) re-certified the benefits of such ensemble methods for enhancing Transformer-based NFs, providing further empirical support for the direction of our research.
Forecast Combinations. Forecast combination is a classic technique for improving forecast accuracy and robustness by leveraging the diverse strengths of multiple models (combiningreview). In this work, we extend this concept (bates1969combination) to the domain of deep learning for long-term time series forecasting. Rather than combining distinct, parallel forecasters into a hybrid model (hybrid), our framework achieves this goal efficiently within a single, structured approach for generating and implicitly combining forecasts within a single neural network. By recursively stacking sub-forecasts, similar in spirit to N-BEATS (nbeats), we ensure that the representation learning is both deep and diverse, allowing the LSTF task to benefit from the full power of modern machine learning.
@@ Appendix C Proof of the approximate proposition
$|\mu_{t}-\frac{1}{c}\sum_{i=1}^{c}{\tilde{y}_{t}^{(i)}}|\leq|\hat{\mu}_{t}-\frac{1}{c}\sum_{i=1}^{c}{\tilde{y}_{t}^{(i)}}|+|\hat{\mu}_{t}-\mu_{t}|\leq|\hat{\mu}_{t}-\frac{1}{c}\sum_{i=1}^{c}{\tilde{y}_{t}^{(i)}}|+b.~~(\text{A3})$ (15)
@@
C.1 Variance Analysis of Total Summed Error
The above proof gives a loose bound as an direct result of the law of large number, which does not depend on any assumption on the temporal dependence. In this part, we consider this factor as a pivot for analysing the effects of different forecasting paradigms, identifying the role of our BDO.
The variance of this total error is:
$var(\sum_{t=1}^{T}e_{t})=\sum_{t=1}^{T}var(e_{t})+2\sum_{t<t^{\prime}}Cov(e_{t},e_{t}^{\prime})$ (18)
For simplicity, let the prediction error at each timestep as $t$ $e_{t}:=\mu_{t}-\frac{1}{c}\sum_{i=1}^{c}{\tilde{y}_{t}^{(i)}}$, and temporarily omit the effect of the predictive bias $b$ of NFM in A3, i.e., $\mu=\hat{\mu}$. Then we can compute the covariance (autocorrelation) of the errors at any two distinct timesteps $k$ and $h$,
[…]
$=\mathbb{E}[\tilde{\mu}_{k}\tilde{\mu}_{h}]-\mu_{k}\mu_{h}$ (20)
$=Cov(\tilde{\mu}_{k}\tilde{\mu}_{h})$ (21)
where we denote $\tilde{\mu}_{t}:=\frac{1}{c}\sum_{i=1}^{c}{\tilde{y}_{t}^{(i)}}$. So we have
$Cov(\tilde{\mu}_{k}\tilde{\mu}_{h})$ $=\frac{1}{c^{2}}\sum_{i=1}^{c}\sum_{j=1}^{c}Cov(y_{k}^{(i)},y_{h}^{(j)})$ (22)
$=\frac{1}{c}Cov(y_{k}^{(i)},y_{h}^{(i)}),~\forall i$ (23)
The last inequality holds by assumption A2, which isolates the correlations across different candidates.
Now we can consider the role of temporal dependence in the error structure.
Case 1: Temporal Independence (The DO Paradigm): The Direct Output (DO) paradigm is designed to satisfy this assumption of temporal independence, as evidenced in Sec. 2. This implies that the error covariance terms in Eq. 23 are zero. Consequently, the total error variance is simply the sum of the per-timestep variances, While this structure effectively prevents the compounding of errors, it often comes at the cost of the model being ”unaware” of the sequential dynamics within the future horizon, potentially leading to higher per-timestep variance. Furthermore, the variance of the total summed error can only benefit from the combination factor $c$ in reducing the individual uncertainty at each future timestep $\sum_{t=1}^{T}var(e_{t})$.
Case 2: Temporal Dependence (The AR Paradigm). : The typical AR paradigm fundamentally violates this assumption and thus leads to errors from one step propagating to the next, which causes the total variance to explode over long horizons. Even though the introduction of the ensembling factor c would reduce the magnitude of this effect, it does not alter the underlying structural problem of error accumulation.
Case 3: A Synthesis (The BDO Paradigm).: Our Boosted Direct Output paradigm operates as a synthesis of these two extremes. It intentionally violates strict temporal independence by recursively feeding sub-forecasts back into the model, thereby making it aware of causal dependencies within the forecast horizon. However, unlike a pure AR model, it would learn a mapping that minimizes the error covariance. Through hierarchical supervision and its patch-wise output structure inherited from DO, the model is trained to control the accumulation of errors. The goal is to leverage the benefits of modeling temporal structure while constraining the error correlation, effectively learning to make the covariance term in Eq. 23 as small as possible.
Note that the above proof and analysis are based on a univariate time series; however, it is easy to extend the result to the multivariate case using a similar process in a certain normed vector space.